The sum of an integer and its reciprocal is 82/9. Find the integer.
Could you please show me how to do this? Thanks so much

Answer by checkley77(12844)

X+1/X=82/9
(X*X+1)/X=82/9
(X^2+1)/X=82/9 NOW CROSS MULTIPLY
9(X^2+1)=82X
9X^2+9-82X=0
9X^2-82X+9=0
(9X-1)(X-9)=0
9X-1=0
9X=-1
X=-1/9 ANSWER.
X-9=0
X=9 ANSWER.